OpenStudy (anonymous):

Short answer. Add the following: (-15x^2+7.9x+8.6)+(-5.5x^2-7.8x+5.2)

OpenStudy (anonymous):

please help me someone with this problem

OpenStudy (saifoo.khan):

Add like terms, \[\Large -20.5 x^2+0.1 x+13.8\]

OpenStudy (anonymous):

-15x^2 + -5.5x^2=-20.5x^2 7.9x+7.8x=0.1x 8.6+5.2=13.8 therefore; -20.5x^2+0.1x+13.8

OpenStudy (anonymous):

so is that the final answer